Three charged particles are arranged on corners of a square. (i) What is the direction of the electric field at the upper right corner?
1
Expert's answer
2020-11-09T10:12:47-0500

Solution

For this question consider side of square is "a" Charges are "q" Figure can be shown as



Electric field on fourth corner is given as

"E_1=\frac{kq}{a^2}"


"E_2=\frac{kq}{a^2}"

Resultant of both above field is

"E'=\frac{\sqrt{2}kq}{a^2}"


"E_3=\frac{kq}{2a^2}"

And finally the resultant electric field will be in direction of E3 on fourth corner is given by

"E=\frac{\sqrt{2}kq}{a^2}+\frac{kq}{2a^2}"


"E=\frac{kq}{a^2}(\sqrt{2}+\frac{1}{2})"

Direction is resultant is 45° from x and y so vector is direction of resultant electric is given by


"E=\frac{kq}{a^2}(\sqrt{2}+\frac{1}{2}) \frac{1}{\sqrt2}(i+j)"